OpenStudy (anonymous):

1. Write the equation of the line with a slope of -5 and a y-intercept of (0,3). 2. Write the equation of the line with a slope of -1/3 and passing through the point (6, -4). 3. Write the equation of the line passing through the points (0, -4) and (-2, 2). 4. Write the equation of the line passing through the points (-6,1) and (-4,2). 5. Write the equation of the line with an undefined slope, passing through the point (2, 5). Hint: refer back to lesson 4.04 for help with this one. 6. Write the equation of the line on the following graph.

OpenStudy (luigi0210):

Word of advice, if you post this many questions, no one will help. Post them one at a time.

OpenStudy (isaiah.feynman):

OK I will do a few, then you can use them as examples to do the rest. Alright??

OpenStudy (anonymous):

works for me

OpenStudy (luigi0210):

@Isaiah.Feynman you're not gonna give answers, right?

OpenStudy (isaiah.feynman):

No boss. Maybe we should split the work I do one example you do another one example, then she can use it to solve the rest.

OpenStudy (luigi0210):

Actually.. I really don't care, as long as you don't get caught giving answers it's fine by me. She should close the question so it's not public and mods don't see it.

OpenStudy (isaiah.feynman):

@bevinlovesyou wanna close the question so we can help without interruptions?

OpenStudy (isaiah.feynman):

@bevinlovesyou the first question is straightforward the equation of a line is y=mx+b, if you know what each term in the equation represents you should easily get your equation.

OpenStudy (anonymous):

that's the problem I don't know anything

OpenStudy (luigi0210):

Well that's a problem

OpenStudy (isaiah.feynman):

Aww. m means slope, b means y-intercept. Now try it out.

OpenStudy (anonymous):

y=5x+(0)(3)

OpenStudy (isaiah.feynman):

Close! but one more thing, since y intercept is the point were the line intersects with the y axis, we pick the y co ordinate of that point, in this case the y co ordinate is 3. Also, the slope is -5 you shouldn't ignore the signs. Try again cutie.

OpenStudy (anonymous):

3=-5x+0?

OpenStudy (anonymous):

@Isaiah.Feynman

OpenStudy (isaiah.feynman):

Actually its y=-5x+3. b is the y intercept, and we use the y co ordinate of this intercept. Thus b = 3. The "y" on the left side is just convention don't bother about it. :)

OpenStudy (anonymous):

ohh okay.....

OpenStudy (anonymous):

still don't get it that much lol

OpenStudy (isaiah.feynman):

I can explain.

OpenStudy (anonymous):

so the y always stays there?

OpenStudy (isaiah.feynman):

Exactly. The only things you worry about is the "mx" and "b".

OpenStudy (isaiah.feynman):

And they are not hard to find.

OpenStudy (anonymous):

can we do # 3 together

OpenStudy (isaiah.feynman):

Sure we can!

OpenStudy (anonymous):

okay walk me through @Isaiah.Feynman

OpenStudy (anonymous):

still there ?

OpenStudy (isaiah.feynman):

Alrighty, so we have two points and they want us to get the equation. Hmm what do we do first?... we find the slope! but the slope is not given :( how to find the slope? *we scratch our heads and come up with a solution* we will use the points given to find the slope! and there is a formula for that too! its \[m =\frac{ y_{2}-y_{1} }{ x_{2}-x_{1}}\] y2 = 2, y1=-4 x2= -2, x1 =0. Okay now can you get the slope so we'll continue?

OpenStudy (anonymous):

Lol your a good teacher keep it coming

OpenStudy (anonymous):

m=-3? @Isaiah.Feynman

OpenStudy (isaiah.feynman):

Yes that's correct!

OpenStudy (anonymous):

your better than my online teacher so it so far will be y=-3x?

OpenStudy (anonymous):

@Isaiah.Feynman

OpenStudy (isaiah.feynman):

we have the slope now, we need b (y intercept). To do this we use the last equation there is in line geometry which is \[y-y_{1} = m(x-x_{1})\]. Now don't fret lol we will ignore the y and x the most useful things there is y1, m, and x1.M is the slope ( we have it already) Y1 and X1 are x and y co ordinates of the points we were given. And guess what? we can choose any point we like!!! so choose the point you like and we will find b!

OpenStudy (anonymous):

Lol I give up again @Isaiah.Feynman

OpenStudy (isaiah.feynman):

No don't give up now we are almost done this question! Simple choose one of the two points given to us in the question. :).

OpenStudy (anonymous):

-4 and -2 @Isaiah.Feynman

OpenStudy (anonymous):

now what ?

OpenStudy (anonymous):

can I have answers instead pweaseee im sleeping :/

OpenStudy (isaiah.feynman):

Umm sorry.. I made a mistake we can't choose any point we like, we must only the first point which is (0,-4). We use this point because its the FIRST point it has x1=0 and y1=-4 and don't forget m=-3!!!. So putting in the values into the equation is showed you we have \[y-(-4)=-3(x-0)\] Simplifying gives us \[y+4=-3x \] Finally, finally we have \[y=-3x-4\] and that's our equation!! Ta da!!!

OpenStudy (anonymous):

yayy thanks <3333

OpenStudy (anonymous):

what about 4? same way

OpenStudy (isaiah.feynman):

Yeah same way!!

OpenStudy (anonymous):

lol will you do it

OpenStudy (isaiah.feynman):

Nope. It would be like repeating question 3 again. :P

OpenStudy (isaiah.feynman):